Telescope and Instrument Status

The following is a list of modifications, problems, or changes to the telescopes and instruments. Please contact to report problems with the telescope or instrumentation.

IRCAM Chip Temperature

Below is a graph of the chip temperature on the IR camera for the past 72 hours, sampled every half hour. 

chip temperature


Nov 6, 2005:  AY122 is finished using IRCAM and Thirty for the Fall.  IRCAM compressor turned off with 5642 hours.  Power to Thirty turned off.  The stopnet script was run on crystalize.

October 27, 2005:  5.0 ultrapure Helium tank down to 900 psi, order more.  Cold head was purged with helium.  Compressor pressure was at 250 psi.  Compressor valve opened to bleed down to 200psi.  Compressor turned back on at 5pm.  During the first hour, the cold head sounded very loud and grindy for a few minutes...this happened a couple of times, but then returns to the thumping sound.

October 26, 2005:  Problem:  IRCAM chip temp was slowly increasing from around 80K to 95K in one week.  In one day the temp increased from 95K to 100K.  Compressor turned off at 8pm to allow warmup.

October 12, 2005:  Problem:  IRCAM chip temp was reaching 107K in the day and staing above 100K at night.
IRCAM was warmed up.  The closed cycle refrigerator lines and head were purged with helium.  The compressor was turned back on around 3pm, @ 5066 hours

September 30, 2005:  Zero corrections found for IRCAM, and pointing looks good for both sides of the meridian.  cra=-0.156 cdec=+0.896.

September 27, 2005:  After finding an alignment star, we set the corrections through tx zero.  Although we can repeat gotos to that star, we can't point to other stars.
For mars, cra=+0.089, cdec=+0.891

September 18, 2005: Replace defective sample & hold (LF398) in premp/levelshift box. Cooler on @ 4494 hours.

The grnding Dec gear has been reworked and the telescope slewing now sounds much better.  The spur gear has been rotated 180 degrees,
cleaned and regreased.  The worm and worm gear have been replaced.  Some bearings have been replaced and the drive system has
been realigned.  The coupler from the drive motor to the telescope shaft has been replaced.  THANK YOU, Terry McDonald and the SSL machine shop!

Leuschner Observatory Index